OpenStudy (anonymous):

Please Help how can you write the expression with a rationalized denominator? 4+3^sqrt 3 / 3^sqrt 6

OpenStudy (anonymous):

@campbell_st @mathmale can you help

OpenStudy (nikato):

Can u draw out the problem?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

That looks confusin^

OpenStudy (anonymous):

\[\frac{ 4+\sqrt[3]{3}}{ \sqrt[3]{6} }\]

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

@nikato

OpenStudy (mathmale):

DB: Thanks for typing this problem into the Equation Editor. SO much easier to read that way. Your task here is to rationalize the den., which means you must eliminate that radical from the denom. Have you done this kind of problem before?

OpenStudy (mathmale):

DB: I'm glad to help you and enjoy helping you, but I do expect to have your full attention during the process. Please respond.

OpenStudy (nikato):

To simplify multiply top and bottom by ur denominator http://www.purplemath.com/modules/radicals7.htm

OpenStudy (mathmale):

DB: Open Study has tattled on you. It says you're "just looking around" (not paying attention to the solution of this problem. DB??

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

DB: your task is to eliminate that radical from the denominator. Your current denominator is 6^(1/3), right?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

6 is not a perfect cube, is it? Our task is to find a multiplier for 6 that transforms it into a perfect cube. Does this sentence make sense to you, and, if so, do you know what to do next?

OpenStudy (anonymous):

we need to find the number clothes to 6 that is a perfect cube

OpenStudy (mathmale):

Not closest to 6, but a number, when multiplied by 6, results in a perfect cube. Sorry, this is a bit abstract and I will have to think for a moment myself to determine what such a number would be. Would you try doing the same thing?

OpenStudy (anonymous):

yes

OpenStudy (nikato):

I found 6^3. Just to speed the process

OpenStudy (mathmale):

So, DB, if you multiply that 6 by 36 (courtesy of nikato), what do you get, and is that result a perfect cube?

OpenStudy (anonymous):

yes its 216

OpenStudy (mathmale):

Good, and is 216 a perfect cube?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Soooo cooool. Now go back to that expression you so beautifully typed in the Equation Editor. Multiply both numerator and den. by 36. Note how the den. becomes a perfect cube?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

So, DB, your denom is 216 and the cube root of that is 6. Your numerator is (4+third root of 3)(36). Write the answer to this problem as cube root of (4+third root of 3)(36)/6. that's it. We're rationalized the denom.

OpenStudy (anonymous):

thanks can you help with one more

OpenStudy (mathmale):

Can't promise; I'm in the midst of helping another student with a different problem and would like to hit the sack soon. But please post your next question; either I or someone else will respond.